# Cheng Lin4, Shiqing Xin5, Yuan Liu3†, Wenping Wang6, Taku Komura1

### v1

@@ 4.2 Comparisons
As a mesh generation framework, faithfully preserving both the overall shape and fine-grained details of the original object is paramount; notable deviations from the reference geometry are unacceptable. To quantitatively measure the fidelity between the generated mesh and the ground-truth shape, we utilize four widely adopted evaluation metrics: Hausdorff Distance (HD), Chamfer Distance (CD), Normal Consistency (NC), and F-score (F1). Following CWF (Xu et al., 2024), we additionally incorporate Edge Chamfer Distance (ECD) and Edge F-score (EF1), as introduced by NMC (Chen & Zhang, 2021), to specifically assess the preservation of sharp features.
@@
Qualitative comparisons in Fig. 9 further illustrate that our method produces meshes with higher fidelity and finer detail. By contrast, MeshAnythingV2 (Chen et al., 2024c) and BPT (Weng et al., 2025) yield meshes of relatively lower quality and resolution, resulting in the loss of high-frequency details. TreeMeshGPT (Lionar et al., 2025) and DeepMesh (Zhao et al., 2025), while capable of generating denser meshes, utilize global one-shot autoregressive mechanisms and consequently struggle to capture complex geometries, such as those evident in the first and last examples. In contrast, our approach leverages a local-to-global prior generation strategy, which not only ensures structural correctness but also enhances the representation of subtle features. Moreover, although datasets like ShapeNet (Chang et al., 2015), Thingi10k (Zhou & Jacobson, 2016), and Objaverse (Deitke et al., 2023b) exhibit varying complexity, our method consistently outperforms competing methods across all of them.
Figure 10: Comparison on triangle numbers.
We present a more compelling example in Fig. 10. For a complex fighter jet model, our method successfully reconstructs intricate details using nearly 30K triangles whereas other approaches struggle with such highly complex shapes, typically yielding only a few hundred to a few thousand triangles. This noticeable gap demonstrates the superior detail recovery and scalability of our method with substantially higher triangle counts.
User Study.
Beyond reconstruction accuracy, it is crucial that generated meshes are artist-quality with sparse, neat, visually compelling, and easy to edit meshes. To assess this, we conducted a user study, sampling 10 models from test datasets. Twenty-seven professional users with expertise in computer graphics or 3D modeling anonymously rated five methods on four criteria: Neatness, Artistry, Similarity to Ground Truth, Detail Recovery. Scores were assigned for the top three methods in each category (3, 2, and 1 points, respectively; 0 for others). The final scores are summarized in Tab. 2. MeshMosaic achieved the highest ratings in all categories, reflecting its superior aesthetic and structural quality. Competing methods scored lower due to issues with mesh stability, single-pass autoregressive models often stall or fail for long, complex meshes, yielding incomplete outputs. BPT (Weng et al., 2025) ranked second in the user study, reflecting similar trends in reconstruction metrics in Tab. 1. Although BPT (Weng et al., 2025) tends to produce more stable outputs, its overall mesh quality is comparatively lower and struggles to preserve fine details. This is further evidenced by its performance in the ECD and EF1 metrics: while BPT (Weng et al., 2025) delivers satisfactory results for relatively simple shapes, such as those in ShapeNet (Chang et al., 2015). Its scores decline markedly with increasing shape complexity (Thingi10K (Zhou & Jacobson, 2016) and Objaverse (Deitke et al., 2023b) datasets).
@@ 5 Discussion
Conclusion.
We present MeshMosaic, a boundary-conditioned local-to-global autoregressive framework that decomposes meshes into compact patches and assembles them coherently. This design fundamentally removes the long-sequence bottleneck and enables higher-resolution quantization, scaling generation to over 100K triangles while preserving fine-grained geometric detail. On ShapeNet (Chang et al., 2015), Thingi10K (Zhou & Jacobson, 2016), and Objaverse (Deitke et al., 2023b) dataset, MeshMosaic achieves state-of-the-art fidelity and perceptual quality, consistently surpassing the baselines. Beyond meshes, it offers a general paradigm for scaling autoregressive generation of structured 3D data via patch-level modeling.
@@ Conclusion.
While MeshMosaic enforces local coherence, boundary conditioning remains primarily local and may leave distant symmetric parts weakly coupled. As illustrated in Fig. 11, the two arms exhibit mild asymmetry despite reasonable connectivity and density. When stronger symmetry is required, this could be alleviated by incorporating global perception mechanisms to couple distant parts. Beyond symmetry, future work may explore cross-patch refinement, multimodal conditioning, and adaptive quantization to further enhance resolution and editability.
@@ A.1 Data Preprocessing
Training datasets are drawn from Objaverse-XL (Deitke et al., 2023a) and other licensed datasets. To enhance data quality, we implemented several filtering procedures. Only meshes with $[500,32000]$ faces are retained, excluding those with excessively low or high token lengths. Meshes are subsequently cleaned and optimized using PyMeshlab (Muntoni & Cignoni, 2021): duplicate vertices/faces removed, closely spaced or overlapping vertices merged, non-manifold elements and edges eliminated.
@@
Fig. 12 provides a comprehensive analysis of our dataset. Moving from left to right, the first graph illustrates the distribution of the number of patches generated through random segmentation. Most simple shapes are divided into fewer than ten patches, whereas a small number of highly complex cases yield over 60 patches. Although our training set contains no more than sixty splits per instance, our method can handle inference tasks involving hundreds of patches during testing. This highlights the strong generalization capability of our method (as shown in Fig. 1).
@@ A.2 Discussion and Ablation
When global point cloud information is omitted (Ours w/o GPC), the network demonstrates reasonable performance in the overfitted regime, as it only needs to reconstruct a single shape. However, in the full dataset setting, the absence of global context leads to significant errors—most notably, the right portion of the mesh exhibits pronounced deformation and collapse, revealing the necessity of global information for guiding overall shape reconstruction. When the GRU-based boundary encoding is eliminated (Ours w/o BD), visible cracks emerge along the seams in both regimes. In addition, the absence of boundary communication induces substantial mesh density asymmetry in the full dataset setting, with adjacent patches developing inconsistencies. This reflects the model’s inability to properly propagate local information between neighboring patches. Disabling the concatenation of boundary tokens for self-attention (Ours w/o SA) again results in prominent seam artifacts, and in the full dataset scenario, produces overlapping, self-intersecting patches. The lack of explicit constraints leads to independent patch generation, which ultimately causes geometric inconsistencies and structural artifacts.
@@
In contrast, the full model employing both boundary and global conditioning produces meshes that are complete, uniform, and visually coherent, with mesh density and topology smoothly balanced across all patches. This clearly demonstrates the effectiveness of our proposed integration of local and global context, and highlights the importance of both conditional mechanisms for high-fidelity mesh generation.
Detail Eecovery.
@@
Although our approach is primarily designed to operate under a segmented training and inference regime, it nevertheless retains the flexibility to infer simple shapes without explicit segmentation. As demonstrated in Fig. 15, both our method and DeepMesh (Zhao et al., 2025) are capable of reconstructing a torus model in the absence of any segmentation.
@@
Further analysis of segmentation strategies is shown in Fig. 16. In the middle example, reconstruction is performed using random segmentation. While the overall shape and fine details can still be recovered, the absence of semantic segmentation often results in patch boundaries that traverse flat or non-essential regions, introducing visual clutter and irregularity into the mesh appearance. By employing PartField (Liu et al., 2025) for semantic guidance, our method achieves noticeably cleaner and more coherent mesh boundaries, significantly enhancing the aesthetic quality without compromising reconstruction fidelity.
@@
Directly scaling DeepMesh (Zhao et al., 2025) to our local-to-global setting is non-trivial. To further demonstrate the benefits of our local-to-global framework and the importance of our boundary condition method, we perform an ablation study that directly compares it with DeepMesh (Zhao et al., 2025). As depicted in Fig. 17, we assess two distinct inference settings for DeepMesh: the first utilizes the entire shape without segmentation, while the second processes each segmented patch individually and subsequently assembles them to form the complete object.
@@
We developed our method on the DeepMesh (Zhao et al., 2025) codebase, thereby ensuring a comparable runtime environment and a rigorous basis for performance assessment. Tab. 3 details the training and inference efficiency of DeepMesh versus our proposed framework, including variants with specific ablations such as the boundary condition encoding (BD) and global point cloud encoding (GPC).
By incorporating GRU-based boundary condition encoding and a global point cloud conditioning module into our pipeline, we necessarily introduce additional computational operations during the training stage. This enhancement results in a moderate increase in training time relative to the original DeepMesh (Zhao et al., 2025) implementation.
@@
[…]
Figure 18: Results generated by MeshMosaic using text prompts (left) or image inputs (right). Initial 3D shapes are created using CLAY (Zhang et al., 2024) and enhanced by our approach.
@@
In Fig. 18, we showcase examples where state-of-the-art SDF-based methods, such as CLAY (Zhang et al., 2024), generate initial 3D geometry from either textual prompts or image inputs. We then refine these preliminary outputs using MeshMosaic, producing artist-quality meshes that retain rich geometric details while optimizing triangle utilization. Compared to the raw outputs from CLAY, meshes processed by our framework exhibit cleaner topology, enhanced visual fidelity, and improved efficiency, making them far better suited for real-world downstream tasks. These results highlight the effectiveness of our method for transforming dense generative outputs into structured, high-quality assets tailored for professional use.
@@
Importantly, despite considerable variations in mesh density, topology, and local connectivity, all generated results exhibit strong global coherence and visual consistency. There are no conspicuous artifacts or discontinuities between regions, confirming that our local-to-global generation strategy supports both creative flexibility and structural integrity across the mesh. This empowers downstream tasks—such as animation, editing, or customization—by facilitating the production of diverse, high-quality assets from unified geometric representations.

### v2

@@ 4.3 Generating single Patch
4.4 Training Strategy
Each mesh is partitioned into patches adaptively in training. Given an input mesh $\mathcal{M}$ with $\mathcal{N}_{p}$ vertices and $\mathcal{N}_{f}$ facets, the number of patches is set to $\mathcal{N}_{\text{seg}}=\frac{\mathcal{N}_{f}}{2000}\times\lambda_{\text{rand}}$, where $\lambda_{\text{rand}}$ is randomly sampled from [0.5, 2.5][0.5,\,2.5] to encourage diversity. The denominator ensures that each patch, after tokenization, yields a sequence length close to the window size ($9\mathrm{K}$) for efficient training. It is worth noting that we do not explicitly constrain the number of patches during inference; the partitioning is determined by the default PartField [liu2025partfield] configuration. Our model flexibly adapts to different patch numbers, handling anywhere from a single patch to several hundred, depending on its geometric detail and richness.
We apply farthest point sampling [moenning2003fast] to select $\mathcal{N}_{\text{seg}}$ points as cluster centers. Voronoi decomposition [aurenhammer1991voronoi] partitions the mesh into patches based on these centers, with bisecting planes separating triangle regions. Clusters are ordered by breadth-first search, starting from the lowest center, to retrieve boundary information sequentially.
We also curate a subset of meshes with high-quality connected component annotations, using each component directly as a patch (with breadth-first ordering, as above). This subset supports tasks requiring more regular, consistent patch segmentation and enables training for semantic reasoning. We provide detailed dataset analysis in Appendix Sec. A.1.
5 Experiments
5.1 Implementation
Our implementation builds upon the released $0.5B$ parameter DeepMesh [zhao2025deepmesh] model, which serves as the base for fine-tuning our approach. We introduce and progressively fuse new boundary conditions and global point cloud features into the architecture, connecting the GRU boundary encoder and global feature input using zero-initialized linear layers. Local point cloud features are mapped directly onto the original input cloud.
The curated dataset consists of 310K meshes, including approximately 90K with connected component information. Please check Appendix Sec. A.1 for the data preprocess and analysis. Training is conducted for seven days on a cluster of 32 NVIDIA H20 96GB GPUs, using a cosine learning rate scheduler that decays from $1\times 10^{-4}$ to $1\times 10^{-5}$. Token window sizes for truncated windows follow DeepMesh’s setting ($9K$, with a $50\%$ overlap). We employ KV-caching in both training and inference and adopt probabilistic sampling (temperature $0.5$) to ensure stable mesh generation.
5.2 Comparisons
To thoroughly assess the effectiveness of our proposed method, we perform comparative experiments with four publicly available state-of-the-art mesh generation methods: MeshAnythingV2 [chen2024meshanythingv2], BPT [bpt], TreeMeshGPT [lionar2025treemeshgpt], and DeepMesh [zhao2025deepmesh]. Note that Hunyuan3D [lei2025hunyuan3dstudio] is a commercial extension of BPT with no available code, making batch evaluation infeasible. For a fair comparison, we therefore use the publicly released BPT-0.5B model, which matches our model size. The comparison includes both quantitative measurements and qualitative visualization. We randomly select 100 samples from each of the ShapeNet [chang2015shapenet], Thingi10K [zhou2016thingi10k], and Objaverse [deitke2023objaverse] datasets for all experiments.
@@ 5.2 Comparisons
As a mesh generation framework, faithfully preserving both the overall shape and fine-grained details of the original object is paramount; notable deviations from the reference geometry are unacceptable. To quantitatively measure the fidelity between the generated mesh and the ground-truth shape, we utilize four widely adopted evaluation metrics: Hausdorff Distance (HD), Chamfer Distance (CD), Normal Consistency (NC), and F-score (F1). Following CWF [xu2024cwf], we additionally incorporate Edge Chamfer Distance (ECD) and Edge F-score (EF1), as introduced by NMC [chen2021neural], to specifically assess the preservation of sharp features.
@@
Qualitative comparisons in Fig. 9 further illustrate that our method produces meshes with higher fidelity and finer detail. By contrast, MeshAnythingV2 [chen2024meshanythingv2] and BPT [bpt] yield meshes of relatively lower quality and resolution, resulting in the loss of high-frequency details. TreeMeshGPT [lionar2025treemeshgpt] and DeepMesh [zhao2025deepmesh], while capable of generating denser meshes, utilize global one-shot autoregressive mechanisms and consequently struggle to capture complex geometries, such as those evident in the first and last examples. In contrast, our approach leverages a local-to-global prior generation strategy, which not only ensures structural correctness but also enhances the representation of subtle features. Moreover, although the three datasets exhibit varying complexity, our method consistently outperforms competing methods across all of them.
We present a more compelling example in the warp figure. For a complex fighter jet model, our method successfully reconstructs intricate details using nearly 30K triangles whereas other approaches struggle with such highly complex shapes, typically yielding only a few hundred to a few thousand triangles. This noticeable gap demonstrates the superior detail recovery and scalability of our method with substantially higher triangle counts.
User Study.
Beyond reconstruction accuracy, it is crucial that generated meshes are artist-quality with sparse, neat, visually compelling, and easy to edit meshes. To assess this, we conducted a user study, sampling 10 models from test datasets. Twenty-seven professional users with expertise in computer graphics or 3D modeling anonymously rated five methods on four criteria: Neatness, Artistry, Similarity to Ground Truth, Detail Recovery. Scores were assigned for the top three methods in each category (3, 2, and 1 points, respectively; 0 for others). The final scores are summarized in Tab. 2. MeshMosaic achieved the highest ratings in all categories, reflecting its superior aesthetic and structural quality. Competing methods scored lower due to issues with mesh stability, single-pass autoregressive models often stall or fail for long, complex meshes, yielding incomplete outputs. BPT [bpt] ranked second in the user study, reflecting similar trends in reconstruction metrics in Tab. 1. Although BPT [bpt] tends to produce more stable outputs, its overall mesh quality is comparatively lower and struggles to preserve fine details. This is further evidenced by its performance in the ECD and EF1 metrics: while BPT [bpt] delivers satisfactory results for relatively simple shapes, such as those in ShapeNet [chang2015shapenet]. Its scores decline markedly with increasing shape complexity (Thingi10K [zhou2016thingi10k] and Objaverse [deitke2023objaverse]).
@@ 6 Discussion
While MeshMosaic enforces local coherence, boundary conditioning remains primarily local and may leave distant symmetric parts weakly coupled. As illustrated in Fig. 10, the two arms exhibit mild asymmetry despite reasonable connectivity and density. When stronger symmetry is required, this could be alleviated by incorporating global perception mechanisms to couple distant parts. Beyond symmetry, future work may investigate multi-node synchronous generation and adaptive quantization to further enhance the speed and quality.
Conclusion.
We present MeshMosaic, a boundary-conditioned local-to-global autoregressive framework that decomposes meshes into compact patches and assembles them coherently. This design fundamentally removes the long-sequence bottleneck and enables higher-resolution quantization, scaling generation to over 100K triangles while preserving fine-grained geometric detail. On ShapeNet [chang2015shapenet], Thingi10K [zhou2016thingi10k], and Objaverse [deitke2023objaverse] dataset, MeshMosaic achieves state-of-the-art fidelity and perceptual quality, consistently surpassing the baselines. Beyond meshes, it offers a general paradigm for scaling autoregressive generation of structured 3D data via patch-level modeling.
@@ A.1 Data Preprocessing
Training datasets are drawn from Objaverse-XL [objaverseXL] and other licensed datasets. To enhance data quality, we implemented several filtering procedures. Only meshes with $[500,32000]$ faces are retained, excluding those with excessively low or high token lengths. Meshes are subsequently cleaned and optimized using PyMeshlab [pymeshlab]: duplicate vertices/faces removed, closely spaced or overlapping vertices merged, non-manifold elements and edges eliminated.
@@
Fig. 11 and Fig. 12 provides a comprehensive analysis of our dataset. Moving from left to right, the first graph illustrates the distribution of the number of patches generated through random segmentation. Most simple shapes are divided into fewer than ten patches, whereas a small number of highly complex cases yield over 60 patches. Although our training set contains no more than sixty splits per instance, our method can handle inference tasks involving hundreds of patches during testing. This highlights the strong generalization capability of our method (as shown in Fig. 1).
@@ A.2 Discussion and Ablation
When global point cloud information is omitted (Ours w/o GPC), the network demonstrates reasonable performance in the overfitted regime, as it only needs to reconstruct a single shape. However, in the full dataset setting, the absence of global context leads to significant errors—most notably, the right portion of the mesh exhibits pronounced deformation and collapse, revealing the necessity of global information for guiding overall shape reconstruction.
When the GRU-based boundary encoding is eliminated (Ours w/o BD), visible cracks emerge along the seams in both regimes. In addition, the absence of boundary communication induces substantial mesh density asymmetry in the full dataset setting, with adjacent patches developing inconsistencies. This reflects the model’s inability to properly propagate local information between neighboring patches.
Disabling the concatenation of boundary tokens for self-attention (Ours w/o SA) again results in prominent seam artifacts, and in the full dataset scenario, produces overlapping, self-intersecting patches. The lack of explicit constraints leads to independent patch generation, which ultimately causes geometric inconsistencies and structural artifacts.
In contrast, the full model employing both boundary and global conditioning produces meshes that are complete, uniform, and visually coherent, with mesh density and topology smoothly balanced across all patches. This clearly demonstrates the effectiveness of our proposed integration of local and global context, and highlights the importance of both conditional mechanisms for high-fidelity mesh generation. It’s worth noting that we did not use PartField [liu2025partfield] for semantic decomposition in this example. Instead, we applied random segmentation to divide the aircraft into three parts. This was done to better visualize the connections along the seams and intentionally make them more noticeable. Our inference with semantic segmentation would not produce seams that are clearly misaligned with the object’s principal orientation.
[…]
@@
Detail Recovery.
@@
Although our approach is primarily designed to operate under a segmented training and inference regime, it nevertheless retains the flexibility to infer simple shapes without explicit segmentation. As demonstrated in Fig. 15, both our method and DeepMesh [zhao2025deepmesh] are capable of reconstructing a torus model in the absence of any segmentation.
[…]
@@
Further analysis of segmentation strategies is shown in Fig. 16. In the middle example, reconstruction is performed using random segmentation. While the overall shape and fine details can still be recovered, the absence of semantic segmentation often results in patch boundaries that traverse flat or non-essential regions, introducing visual clutter and irregularity into the mesh appearance. By employing PartField [liu2025partfield] for semantic guidance, our method achieves noticeably cleaner and more coherent mesh boundaries, significantly enhancing the aesthetic quality without compromising reconstruction fidelity.
@@
Directly scaling DeepMesh [zhao2025deepmesh] to our local-to-global setting is non-trivial. To further demonstrate the benefits of our local-to-global framework and the importance of our boundary condition method, we perform an ablation study that directly compares it with DeepMesh [zhao2025deepmesh]. As depicted in Fig. 17, we assess two distinct inference settings for DeepMesh: the first utilizes the entire shape without segmentation, while the second processes each segmented patch individually and subsequently assembles them to form the complete object.
@@
[…]
Figure 18: Results generated by MeshMosaic using text prompts (left) or image inputs (right). Initial 3D shapes are created using CLAY [CLAY] and enhanced by our approach.
[…]
Figure 19: Diversity of our generation results.
[…]
Figure 20: Gallery of our artist mesh generation results.
@@
We developed our method on the DeepMesh [zhao2025deepmesh] codebase, thereby ensuring a comparable runtime environment and a rigorous basis for performance assessment. Tab. 3 details the training and inference efficiency of DeepMesh versus our proposed framework, including variants with specific ablations such as the boundary condition encoding (BD) and global point cloud encoding (GPC).
By incorporating GRU-based boundary condition encoding and a global point cloud conditioning module into our pipeline, we necessarily introduce additional computational operations during the training stage. This enhancement results in a moderate increase in training time relative to the original DeepMesh [zhao2025deepmesh] implementation.
@@
In Fig. 18, we showcase examples where state-of-the-art SDF-based methods, such as CLAY [CLAY], generate initial 3D geometry from either textual prompts or image inputs. We then refine these preliminary outputs using MeshMosaic, producing artist-quality meshes that retain rich geometric details while optimizing triangle utilization. Compared to the raw outputs from CLAY, meshes processed by our framework exhibit cleaner topology, enhanced visual fidelity, and improved efficiency, making them far better suited for real-world downstream tasks. These results highlight the effectiveness of our method for transforming dense generative outputs into structured, high-quality assets tailored for professional use.
@@
Importantly, despite considerable variations in mesh density, topology, and local connectivity, all generated results exhibit strong global coherence and visual consistency. There are no conspicuous artifacts or discontinuities between regions, confirming that our local-to-global generation strategy supports both creative flexibility and structural integrity across the mesh. This capability enables downstream tasks such as animation, editing, and customization by supporting the generation of diverse, high-quality assets from a unified geometric representation.
More Results
Finally, we present more of our results in Fig. 20 to demonstrate the powerful capabilities of MeshMosaic.